OpenStudy (anonymous):

which number is divisible by 2,3,4, and 6. A. 124. B.3.072. C.1,005. D.786.

OpenStudy (lieutenantgeneral):

get a calculator and divide them by the answers, which ever one equals a decimal is the wrong answer

OpenStudy (anonymous):

its pretty easy when you think about it

OpenStudy (anonymous):

ok thanks

OpenStudy (lieutenantgeneral):

i did this before

OpenStudy (lieutenantgeneral):

welcome

OpenStudy (anonymous):

12 From Mathematica: FactorInteger[{2, 3, 4, 6}] -> {{{2, 1}}, {{3, 1}}, {{2, 2}}, {{2, 1}, {3, 1}}}

OpenStudy (anonymous):

\[\frac{1}{12} \{124,3072,1005,786\}\to \left\{\frac{31}{3},256,\frac{335}{4},\frac{131}{2}\right\} \]

OpenStudy (anonymous):

B.

OpenStudy (campbell_st):

there are a number of simple divisiblity rules you can use divisible by 2 its an even number, so that rules out C

OpenStudy (cwrw238):

just try and divide each number by 12 3072 is the only one one

OpenStudy (campbell_st):

divisible by 3, the digits add to a multiple of 3 1 + 2 + 4 = 7 so that eliminates A

OpenStudy (campbell_st):

divisible by 4, the last 2 digits are a multiple of 4.... 86/4 = 21.5 so that eliminates D

OpenStudy (campbell_st):

divisible by 6, its even and the digits add to a multiple of 3. so 3 + 0 + 7 + 2 = 12 its even the last 2 digits 72 = 4 x 18 so it meets all the conditions

OpenStudy (campbell_st):

just remember some simple divisibility rules... ends in 5 or 0.... divisible by 5 ends in 0 divisible by 10